# Simple Calibration via Geodesic Kernels

### v9

@@ 4.2 Conclusions
In this paper, we demonstrate a simple intuition that transforms traditional deep discriminative models into a type of binning and kerneling approach. The bin boundaries are determined by the internal structure learned by the parent approach and Geodesic distance encodes the low-dimensional structure learned by the model. Moreover, geodesic distance introduced in this paper may have broader impact on understanding the internal structure of the deep discriminative models, which may be of interest for future work. Our code, including the package and experiments in this manuscript, will be made publicly available upon acceptance of the paper. Appendix Section A answers a number of frequently asked questions about this work.
@@ Acknowledgements
The authors thank the support of the NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (NSF grant 2031985) and THEORINET. This work is graciously supported by the Defense Advanced Research Projects Agency (DARPA) Lifelong Learning Machines program through contracts FA8650-18-2-7834 and HR0011-18-2-0025. Research was partially supported by funding from Microsoft Research and the Kavli Neuroscience Discovery Institute. We also thank the helpful discussion with Will LeVine, Tyler M. Tomita, Ali Geisa, Tiffany Chu and Jacob Desman.
@@ Appendix A Frequently Asked Questions
Many people were concerned about the runtime of our approach, possibly it could be an exponential function of the number of nodes. However, we did additional experiments (see Fig. 6) where we show training and testing time both are linear in the number of the nodes. Moreover, training time is only 200 seconds even when there are 40,000 nodes running on a MacBook Pro with an Apple M1 Max chip and 64 GB of RAM. The number of total polytopes in KDN is upper bounded by the training sample size as we only consider the polytopes populated by training data (see the first paragraph of Section 2.3 and Equation 7).
Figure 6: Scaling plots for KDN on datasets with different feature dimensions. The training sample size is fixed at $10$,$000$ and the number of nodes in each layer is increased in a four layer fully connected network. Total number of polytopes populated by the training data saturates to training size (left) while training (middle) as well as testing time (right) scales linearly with node number.

### v10

@@ 4.2 Conclusions
In this paper, we have demonstrated a simple intuition that transforms traditional deep discriminative models into a type of binning and kerneling approach. The bin boundaries are determined by the internal structure learned by the parent approach, and the Geodesic distance encodes the low-dimensional structure learned by the model. The vision experiments in this paper show how the model can be applied to deeper networks by employing a front-end encoder to extract local image features (as explained in the first paragraph of Section 3.2). Instead of applying KDN to the entire ViT, we applied it solely to the final fully connected layers with ReLU activations. This reduces the effective depth of the model, ensuring that KDN can be applied even to encoders that do not rely on ReLU activations. Moreover, the geodesic distance introduced in this paper may have a wider impact on understanding the internal structure of the deep discriminative models, which may be of interest for future work. However, a key limitation of our approach is the need to store parameters for each populated polytope, which may hinder scalability to very large datasets. One potential solution is to merge nearby polytopes into a single representative polytope. Our code, including the package and the approach proposed in this manuscript, is available from https://github.com/neurodata/kdg. Appendix Section A answers a number of frequently asked questions about this work.
@@ Acknowledgements
The authors thank the support of the NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (NSF grant 2031985) and THEORINET. This work was graciously supported by the Defense Advanced Research Projects Agency (DARPA) Lifelong Learning Machines program through contracts FA8650-18-2-7834 and HR0011-18-2-0025. Research was partially supported by funding from Microsoft Research and the Kavli Neuroscience Discovery Institute. We also thank the helpful discussion with Will LeVine, Tyler M. Tomita, Ali Geisa, Tiffany Chu and Jacob Desman.
@@ Appendix A Frequently Asked Questions
Many people were concerned about the runtime of our approach, possibly it could be an exponential function of the number of nodes. However, we did additional experiments (see Fig. 6) where we show training and testing time both are linear in the number of the nodes. Moreover, training time is only 200 seconds even when there are 40,000 nodes running on a MacBook Pro with an Apple M1 Max chip and 64 GB of RAM. The number of total polytopes in KDN is upper bounded by the training sample size as we only consider the polytopes populated by training data (see the first paragraph of Section 2.3 and Equation 7). With overparameterized networks, which is the norm, they typically exhibit this property as shown in Fig. 6 that the number of populated polytopes is equal to the sample size
Figure 6: Scaling plots for KDN on datasets with different feature dimensions. The training sample size is fixed at $10$,$000$ and the number of nodes in each layer is increased in a four layer fully connected network. Total number of polytopes populated by the training data saturates to training size (left) while training (middle) as well as testing time (right) scales linearly with node number. With over-parameterized networks, which is the norm, they typically exhibit this property that the number of populated polytopes is equal to the sample size.